# DiverseNet: Decision Diversified Semi-supervised Semantic Segmentation Networks for Remote Sensing Imagery

### v2

@@ 6.2 Quantitative Results and Analysis
Since the DiverseModel technique uses three different segmentation networks, we also evaluated the performance of each member network within the DiverseModel on the Potsdam dataset in Table 6. The component models UNet, SegNet, and PSPNet use unlabeled data through the way of SHS. DiverseModel demonstrates a significant improvement across all performance metrics in the segmentation task compared to each network. In particular, the metric of PA experiences a notable improvement of 4.57% compared to the best-performing individual component model. This phenomenon can be attributed to the enhancement of pseudo-label diversity through cross-model supervision, resulting in a significant improvement in PA (recall) during the test phase. The findings suggest that the cross-supervision of different networks has the potential to achieve superior performance compared to the best-performing individual component.
Potsdam DFC2020 RoadNet Massachusetts
Image [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Ground Truth [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
MT [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
IoU = 39.38% IoU = 70.08% IoU = 79.30% IoU = 73.37%
CCT [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
IoU = 51.19% IoU = 78.21% IoU = 86.24% IoU = 75.97%
GCT [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
IoU = 40.35% IoU = 76.66% IoU = 80.96 IoU = 74.20%
Potsdam DFC2020 RoadNet Massachusetts
CPS Refer to caption Refer to caption Refer to caption Refer to caption
IoU = 50.21% IoU = 79.39% IoU = 84.92% IoU = 74.16%
DiverseModel Refer to caption Refer to caption Refer to caption Refer to caption
IoU = 58.76% IoU = 80.28% IoU = 94.67% IoU = 76.84%
DiverseHead (Dropout) Refer to caption Refer to caption Refer to caption Refer to caption
IoU = 57.46% IoU = 80.12% IoU = 93.01% IoU = 78.93%
DiverseHead (Dynamic Freezing) Refer to caption Refer to caption Refer to caption Refer to caption
IoU = 58.25% IoU = 80.31% IoU = 90.35% IoU = 78.17%
Figure 7: Visual results of each method on four datasets. The IoU score of prediction for each method is shown in its sub-caption.
6.3 Qualitative Results and Analysis
@@
The visible results are presented in Figures 7. For each dataset, we randomly selected one case of the RGB image, its ground truth, and predictions for all the methods. Also, the IoU scores are calculated for all predictions with the ground truth for each case, which are shown in the sub-caption of each prediction. Based on the IoU values of visual predictions, the highest score is obtained by either DiverseModel or DiverseHead (DT&DF). Especially for the RoadNet dataset, both the DiverseModel and DiverseHead families achieve an IoU of over 90% surpassing other methods by at least 3.5%. Visually, the segmentation maps of DiverseModel and DiverseHead display better continuity and show better similarity to the ground truth than other methods.
7 Conclusion
@@
In this paper, we proposed a lightweight semi-supervised learning approach based on a multi-head structure called DiverseHead. Based on the multi-head structure we provide two perturbation methods, namely dynamic freezing and dropout. Taking inspiration from the theory of bagging, a voting mechanism is proposed to generate beneficial pseudo-labels in the training stage. This simple and lightweight semi-supervised learning framework shows competitive performance for the segmentation of remote sensing imagery. Furthermore, we conducted further analysis and additional evaluation on the previously proposed multi-network-based semi-supervised learning method known as DiverseModel. Based on the results obtained from the aforementioned four remote sensing datasets, DiverseHead and DiverseModel demonstrate comparable performance while significantly outperforming various widely used semi-supervised learning frameworks. It is important to highlight that the training structure of DiverseHead is significantly lighter than that of DiverseModel and other state-of-the-art methods, while still achieving competitive performance. This implies that DiverseHead is a better option for those who lack high-memory computation resources.

### v3

@@ VI-B Quantitative Results and Analysis
VI-C Qualitative Results and Analysis
@@
The visible results are presented in Figures 9. For each dataset, we randomly selected 2 cases of the original image, its ground truth, and predictions for the proposed methods and various classic methods. Also, the mIoU scores are calculated for all predictions with the ground truth for each case, which are shown in the sub-caption of each prediction. Based on the IoU values of visual predictions, the highest score is obtained by either DiverseModel or DiverseHead (DT&DF) for most cases. Especially for cases of the RoadNet dataset, both the DiverseModel and DiverseHead families achieve a mIoU of over 90% surpassing other methods by at least 6.35%. Visually, the segmentation maps of DiverseModel and DiverseHead display better overall similarity to the ground truth than other methods.
Potsdam DFC2020 RoadNet Massa- chusetts Potsdam DFC2020 RoadNet Massa- chusetts
(1) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
(2) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
(3) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
64.08% 80.99% 81.40% 76.12% 41.88% 70.08% 85.33% 75.90%
(4) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
69.24% 85.86% 85.44% 78.03% 47.18% 78.21% 86.93% 77.44%
(5) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
68.03% 82.38% 81.57% 76.56% 47.41% 76.66% 83.61% 72.72%
(6) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
67.19% 87.44% 85.14% 77.48% 45.06% 79.39% 86.21% 77.09%
(7) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
74.81% 88.06% 94.67% 79.10% 51.82% 80.28% 92.95% 78.71%
(8) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
73.39% 88.32% 94.13% 81.63% 50.43% 80.12% 90.83% 78.89%
(9) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
76.13% 88.30% 91.79% 80.34% 55.21% 80.31% 92.04 80.87%
Figure 9: Visual results of each method on four datasets. The mIoU score of the prediction for each method is shown in its sub-caption. The numbers in the first column are – (1): Image (2): Ground Truth (3): MT (4): CCT (5): GCT (6): CPS (7): DiverseModel (8): DiverseHead(DT) (9): DiverseHead(DF)
VII Conclusion
@@
In this paper, we proposed a lightweight and efficient semi-supervised learning approach based on a multi-head structure called DiverseHead. Based on the multi-head structure, we provide two perturbation methods, namely dynamic freezing and dropout. Taking inspiration from the theory of bagging, a voting mechanism is proposed to generate beneficial pseudo-labels in the training stage. This simple and lightweight semi-supervised learning framework shows competitive performance for the segmentation of remote sensing imagery. Also, It can be easily combined with state-of-the-art SSL methods and further improve their performance. Furthermore, we conducted further analysis and additional evaluation on the previously proposed multi-network-based semi-supervised learning method known as DiverseModel. Based on the results obtained from the aforementioned four remote sensing datasets, DiverseHead and DiverseModel demonstrate comparable performance while significantly outperforming various classic semi-supervised learning frameworks. From the application perspective, the proposed DiverseNet could theoretically be utilised for a wide range of image-based tasks, including medical imaging, remote sensing, and natural image segmentation. Having evaluated the methods on datasets with multi-band images, our approach is likely to perform competitively in MRI and other multi-band image segmentation tasks. We will explore more applications using the proposed method in the future. For technical improvement, future work could explore investigating the integration of the two proposed and other perturbation strategies in SSL, and designing more sophisticated interaction mechanisms among component models and optimising the hyperparameters to enhance their synergy.